OpenStudy (bossimbacon):

Which shows this equation in standard form? 6 - 2(x - y) = 3x + 5 A. 5x - 2y = 1 B. 5x - 2y = -1 C. 5x + 2y = 1 D. 2x - 5y = 1 @StudyGurl14

OpenStudy (anonymous):

its B since 6-2(x-y)=3x+5 6-2x+2y=3x+5 6-5+2y=3x+2x 1+2y=5x 1=5x-2y